# Interpretable Image Emotion Recognition: A Domain Adaptation Approach Using Facial Expressions

### v3

@@ Abstract
This paper proposes a feature-based domain adaptation technique for identifying emotions in generic images, encompassing both facial and non-facial objects, as well as non-human components. This approach addresses the challenge of the limited availability of pre-trained models and well-annotated datasets for Image Emotion Recognition (IER). Initially, a deep-learning-based Facial Expression Recognition (FER) system is developed, classifying facial images into discrete emotion classes. Maintaining the same network architecture, this FER system is then adapted to recognize emotions in generic images through the application of discrepancy loss, enabling the model to effectively learn IER features while classifying emotions into categories such as ‘happy,’ ‘sad,’ ‘hate,’ and ‘anger.’ Additionally, a novel interpretability method, Divide and Conquer based Shap (DnCShap), is introduced to elucidate the visual features most relevant for emotion recognition. The proposed IER system demonstrated emotion classification accuracies of 60.98% for the IAPSa dataset, 58.86% for the ArtPhoto dataset, 69.13% for the FI dataset, and 58.06% for the EMOTIC dataset. The system effectively identifies the important visual features leading to specific emotion classifications and provides detailed embedding plots to explain the predictions, enhancing the understanding and trust in AI-driven emotion recognition systems.
@@ 1 Introduction
Humans predominantly portray emotion-related information visually. The images and videos portraying various emotions may contain facial contents, human activities, different backgrounds, and non-human objects. There is a need to develop computational systems that can identify the emotional information expressed in them. These systems study people’s emotional responses to visual information [19], and they find their usage in animation, gaming, marketing, entertainment, graphics, and lie detection [23, 19]. The semantic-level Image Emotion Recognition (IER) analysis [40] has been explored by the researchers; however, affective level analysis is more difficult [23, 51, 42]. The recent progress of deep learning has caused an enormous performance boost for image recognition and object classification [25, 40]. Deep learning-based approaches have been successfully used for Facial Expression Recognition (FER); however, emotion analysis in generic images is a very complex task as they may include human, non-facial components, and non-human objects [27]. The emotional expression in images is sometimes attributed to high-level visual features such as background, and facial structure, whereas, sometimes, low-level image features such as texture, edge, color, and shape determine the emotional portrayal in images.
@@
In the context of emotion recognition in the visual domain, while FER has been extensively studied [36, 45, 56, 59], its methodologies cannot be directly transferred to IER for generic images without adaptation. The absence of sufficient pre-trained models and well-annotated datasets for IER compared to FER presents significant challenges. To address this gap, the proposed system uses domain adaptation to leverage the advances in FER to enhance IER. Domain adaptation is a special case of transfer learning that is applicable when the source and target domain share identical feature spaces but have different data distributions. The inspiration behind utilizing FER advancements for IER stems from the understanding that emotions, whether captured through facial expressions or generic images, share underlying affective properties that can be modelled similarly [47]. This approach promises to bridge the gap in current IER technologies by adapting successful FER methodologies to broader emotional contexts in images.
This paper proposes an interesting problem: how to address image emotion recognition using facial expression recognition. The basic idea is to constrain the output probabilities of these two tasks to be consistent with each other. This is achieved through a discrepancy loss that is applied across the two cross-dataset tasks, incorporating a comprehensive architecture combining three modules together, thus resulting in a reasonably good interpretable system for emotion recognition [40, 51].
@@
The proposed system classifies an image into discrete categories using an adapted FER model, constructed using VGG16, residual blocks, convolution, max pooling, and dense layers, trained with the Adam optimizer. This main FER model and the adapted IER model are trained simultaneously. The discrepancy loss minimizes differences between the models, enabling the adapted IER model to effectively learn the IER domain’s distribution. The DnCShap interpretability approach, detailed later in Section 3.4.1, enhances this model by identifying the key features contributing most towards recognizing specific emotion classes.
Domain adaptation is central to our approach, focusing specifically on feature-based strategies. By maintaining the same architecture for both FER and IER and employing discrepancy loss, we align the target domain’s distribution without altering the network architecture, allowing for seamless adaptation across both domains [46, 54, 61]. This feature-based strategy enhances the model’s ability to generalize across facial and generic images, ensuring that our system effectively learns IER features while maintaining robustness in FER applications.
@@
The experimental results on the International Affective Picture System Subset a (IAPSa), ArtPhoto, Flickr & Instagram, and The Emotions In Context (EMOTIC) datasets demonstrate emotion recognition accuracies of 60.98%, 58.86%, 69.13% and 58.06% respectively, which significantly exceed those reported in previous studies, validating the effectiveness of our approach. The architecture of the proposed IER system and the corresponding domain adaptation strategy have been refined through these studies, highlighting the critical link between the identified gaps in existing models and our contributions to addressing these challenges. The key contributions of this paper are listed as follows.
@@ 2.1.1 Facial Expression Recognition
The FER has significantly evolved from traditional manual feature extraction methods to advanced deep learning approaches [16, 45, 3]. Traditional FER involves a multi-step process starting with pre-processing, where techniques like histogram equalization are commonly applied to normalize lighting across images [16, 67]. This is followed by feature extraction using methods such as local binary patterns, which capture textural information [9], and Gabor wavelets, noted for their edge detection capabilities [2]. These traditional techniques, however, often struggle with the variability of facial expressions within the same category. Transitioning to deep learning, these methods have revolutionized FER by enhancing the ability to handle intra-class variations effectively, without the need for manual intervention [68]. Khorrami et al. demonstrated this with their work on a deep network with zero bias, which showed improved performance on the Cohn-Kanade (CK) dataset [22]. Further advancements include Aneja et al.’s deep FER model that facilitates animated and human facial expression mapping [1], highlighting the versatility of deep learning models.
@@ 2.2 Domain Adaptation for Image Emotion Recognition
Historically, emotion recognition has relied heavily on facial images [45], with systems often utilizing attention models to enhance FER [13]. Domain adaptation in this field encompasses various approaches including ‘feature-based adaptation’ where features are transformed or augmented to bridge the gap between source and target domains [55, 37], ‘instance-based adaptation’ which focuses on selecting relevant training samples from the source domain [58], and ‘model-based adaptation’ strategies that involve modifying the model architecture or training strategy, such as using adversarial techniques or discrepancy loss [14, 54].
We employ a feature-based domain adaptation approach, focusing directly on visual modalities to preserve the integrity of emotional data, thereby avoiding the information loss associated with methods like those using image captions for domain adaptation [32]. This approach is exemplified by the FaceNet architecture, which learns mappings from facial images into representational spaces based on face-similarity measures [56]. However, these FER models often falter when applied to non-facial images, necessitating adaptations for broader IER applications. Our system adapts pre-trained FER models to effectively recognize emotions in generic non-facial images using this feature-based domain adaptation strategy. Employing discrepancy loss to align the feature distributions of FER and IER models without altering the network architecture ensures optimal transfer and adaptation of learned features across domains [54, 61]. This method addresses the challenge of varied emotional cues in different image contexts and is capable of handling diverse datasets. It not only enhances the accuracy and applicability of our system but also interprets a wide range of visual features, significantly improving the robustness of emotion recognition.
@@ 2.3 Interpretability and Explainability
Interpretability and Explainability are increasingly vital in the field of AI and machine learning, particularly within the domain of emotion recognition [33, 66, 31]. These concepts are essential for building trust in AI systems, as they allow users to understand and validate the decision-making processes of the models [4]. Interpretable emotion analysis has become a significant trend, with research expanding across both visual [64, 48] and other modalities [38, 29], underscoring the need for systems that users can understand and interact with confidently.
@@
Our research builds on these foundational works by proposing a system that not only explains emotion recognition outcomes but also provides insights into the training dynamics of network weights. This is crucial for the accurate interpretation of subtle emotional cues, which is essential in effective human-computer interaction. By enhancing transparency and interpretability, we aim to create more reliable AI tools for emotion analysis [30]. Our findings show that integrating real-time data interpretation and dynamic adjustment of network weights not only meets but exceeds theoretical expectations and outperforms traditional models in initial tests. These results underline the practical benefits of our approach in addressing gaps in existing models effectively.
@@ 3.3 Image Emotion Recognition
Domain Adaptation Strategy: We employ a feature-based domain adaptation strategy using the same network architecture for both FER and IER tasks. This approach utilizes discrepancy loss to align the target domain’s distribution with that of the source effectively. The discrepancy loss, which is minimized after each training epoch, helps the adapted IER model accurately learn the IER domain’s distribution, despite the differing marginal distributions of data points between the FER and IER domains. This alignment is critical for the model’s ability to perform consistently across both domains. The overall loss function is detailed in Eq. 1.
@@ 4.1 Datasets
FER Datasets: The following FER datasets have been used.
  • FER13 [6]: This dataset includes 35,887 emotion-labeled images with variations such as facial occlusions, and incomplete faces.
  • FERG [1]: It consists of 55,767 annotated and animated faces from six characters, along with emotion labels.
  • JAFFE [20]: It includes 213 images of Japanese female models portrayed along with different facial expressions.
  • CK+ [41]: This dataset contains 123 objects’ 927 images portrayed along with different emotions.
IER Datasets: The following IER datasets have been used.
  • IAPSa dataset [44] - The subset A of the IAPS dataset, i.e., IAPSa dataset, has been prepared by Mikels et al. [44] by taking 395 images and mapping their arousal and valence values to discrete emotion categories.
  • ArtPhoto dataset [42] - It contains 806 art photos whose emotions are labelled by the artists who prepared the photos.
  • Flicker & Instagram (FI) dataset [73] - It contains 23,308 images collected by You et al. [73] from Flicker and Instagram and labelled with discrete emotion categories.
  • EMOTIC dataset [24] - It contains images involving people in various backgrounds and contexts. The images are labelled with continuous dimensions and discrete categories.
4.2 Training Strategy and Hyperparameter Setting

### v4

@@ Abstract
This paper proposes a feature-based domain adaptation technique for identifying emotions in generic images, encompassing both facial and non-facial objects, as well as non-human components. This approach addresses the challenge of the limited availability of pre-trained models and well-annotated datasets for Image Emotion Recognition (IER). Initially, a deep-learning-based Facial Expression Recognition (FER) system is developed, classifying facial images into discrete emotion classes. Maintaining the same network architecture, this FER system is then adapted to recognize emotions in generic images through the application of discrepancy loss, enabling the model to effectively learn IER features while classifying emotions into categories such as ‘happy,’ ‘sad,’ ‘hate,’ and ‘anger.’ Additionally, a novel interpretability method, Divide and Conquer based Shap (DnCShap), is introduced to elucidate the visual features most relevant for emotion recognition. The proposed IER system demonstrated emotion classification accuracies of 61.86% for the IAPSa dataset, 62.47% for the ArtPhoto dataset, 70.78% for the FI dataset, and 59.72% for the EMOTIC dataset. The system effectively identifies the important visual features that lead to specific emotion classifications and also provides detailed embedding plots explaining the predictions, enhancing the understanding and trust in AI-driven emotion recognition systems.
@@ 1 Introduction
Humans predominantly portray emotion-related information visually. The images and videos depicting various emotions may contain facial contents, human activities, different backgrounds, and non-human objects. There is a need to develop computational systems that can identify the emotional information expressed in them. These systems study people’s emotional responses to visual information [22] and are used in animation, gaming, marketing, entertainment, graphics, and lie detection [27, 22]. The semantic-level Image Emotion Recognition (IER) analysis [47] has been explored by the researchers; however, affective level analysis is more difficult [27, 60, 50]. The recent progress of deep learning has caused an enormous performance boost for image recognition and object classification [29, 47]. Deep learning-based approaches have been successfully used for Facial Expression Recognition (FER); however, emotion analysis in generic images is a very complex task as they may include human, non-facial components, and non-human objects [31]. The emotional expression in images is sometimes attributed to high-level visual features such as background, and facial structure, whereas, sometimes, low-level image features such as texture, edge, color, and shape determine the emotional portrayal in images.
@@
In the context of emotion recognition in the visual domain, while Facial Expression Recognition (FER) has been extensively studied [40, 53, 65], its methodologies cannot be directly transferred to Image Emotion Recognition (IER) for generic images without adaptation. The absence of sufficient pre-trained models and well-annotated datasets for IER compared to FER presents significant challenges. To bridge this gap, our proposed system employs a domain adaptation strategy, a special case of transfer learning where the source and target domains share identical feature spaces but have different data distributions. This approach utilizes the advances in FER to enhance IER by adapting successful FER methodologies to broader emotional contexts in images, ensuring the system effectively learns IER features while maintaining robustness in FER applications. By maintaining the same architecture for both FER and IER and employing discrepancy loss, we align the target domain’s distribution without altering the network architecture, allowing for seamless adaptation across both domains [54, 63, 56]. This feature-based strategy enhances the model’s ability to generalize across facial and generic images, which is inspired by the understanding that emotions, whether captured through facial expressions or generic images, share underlying affective properties that can be modeled similarly.
@@
The proposed system classifies an image into discrete categories using an adapted FER model, constructed using VGG16, residual blocks, convolution, max pooling, and dense layers, trained with the Adam optimizer. This main FER model and the adapted IER model are trained simultaneously. The discrepancy loss minimizes differences between the models, enabling the adapted IER model to effectively learn the IER domain’s distribution. The DnCShap interpretability approach, detailed later in Section 3.4.1, enhances this model by identifying the key features contributing most towards recognizing specific emotion classes. While deep learning methods have achieved good performance in IER, they often function as “black-box” solutions with limited transparency [37]. In the context of IER, interpretability is particularly crucial because emotional cues can be subjective and multifaceted, making it essential for practitioners and end-users to understand how decisions are being made. This lack of interpretability poses challenges in real-world applications where accountability and user trust are paramount [35, 7]. The proposed DnCShap interpretability approach addresses these concerns by highlighting the salient features that most strongly influence model predictions, thereby bridging the gap between high accuracy and user-centric transparency in IER systems. By revealing the internal decision-making pathways, DnCShap ensures that stakeholders can comprehend the rationale behind specific emotion classifications, ultimately enhancing trust and facilitating broader adoption in sensitive or safety-critical domains.
The proposed domain adaptation approach minimizes discrepancy loss, enabling an effective transfer of learned representations from FER to broader IER task. This is achieved by systematically aligning source (facial) and target (generic) distributions and constraining the output probabilities of these two tasks to be consistent through the use of a discrepancy loss. Compared to existing methods, which typically tackle face-only or highly constrained image data, the proposed system preserves robust features pertinent to emotion classification. Furthermore, the newly introduced DnCShap interpretability module provides explanations at both feature and layer-wise levels, detailing what parts of an image drive the model’s decisions—an aspect rarely addressed in prior domain adaptation setups. By uniting robust domain adaptation with detailed interpretability, this results in an interpretable emotion recognition system, advancing the state-of-the-art in IER [47, 60].
@@
The experimental results on the International Affective Picture System Subset a (IAPSa), ArtPhoto, Flicker & Instagram, and The Emotions In Context (EMOTIC) datasets demonstrate emotion recognition accuracies of 61.86%, 62.47%, 70.78% and 59.72% respectively, which significantly exceed those reported in previous studies, validating the effectiveness of our approach. The architecture of the proposed IER system and the corresponding domain adaptation strategy have been refined through these studies, highlighting the critical link between the identified gaps in existing models and our contributions to addressing these challenges. The key contributions of this paper are listed as follows.
@@ 2.1.1 Facial Expression Recognition
The FER has significantly evolved from traditional manual feature extraction methods to advanced deep learning approaches [19, 53, 6]. Traditional FER involves a multi-step process starting with pre-processing, where techniques like histogram equalization are commonly applied to normalize lighting across images [19]. This is followed by feature extraction using methods such as local binary patterns, which capture textural information [12], and Gabor wavelets, noted for their edge detection capabilities [5]. These traditional techniques, however, often struggle with the variability of facial expressions within the same category. Transitioning to deep learning, these methods have revolutionized FER by enhancing the ability to handle intra-class variations effectively, without the need for manual intervention [75]. Khorrami et al. demonstrated this with their work on a deep network with zero bias, which showed improved performance on the Cohn-Kanade (CK) dataset [26]. Further advancements include Aneja et al.’s deep FER model that facilitates animated and human facial expression mapping [3], highlighting the versatility of deep learning models.
@@ 2.2 Domain Adaptation for Image Emotion Recognition
Historically, emotion recognition has relied heavily on facial images [53], with systems often utilizing attention models to enhance FER [17]. Domain adaptation in this field encompasses various approaches including ‘feature-based adaptation’ where features are transformed or augmented to bridge the gap between source and target domains [64, 41], ‘instance-based adaptation’ which focuses on selecting relevant training samples from the source domain [68], and ‘model-based adaptation’ strategies that involve modifying the model architecture or training strategy, such as using adversarial techniques or discrepancy loss [63].
Recent works have extended domain adaptation into cross-subject and multi-modal directions specifically targeting emotion recognition. Lin et al. [44] introduced a multi-source domain adaptation framework for visual sentiment classification, demonstrating how fusing knowledge across multiple labeled domains can significantly enhance transferability. Zhao et al. [83] proposed a “plug-and-play” approach to domain adaptation for EEG-based emotion recognition, highlighting the modular reusability of adaptation blocks across different subjects. Similarly, Ahn et al. [1] investigated cross-corpus speech emotion recognition through few-shot learning and domain adaptation, underscoring the interplay between minimal supervision and acoustic domain shifts. Meanwhile, Zheng et al. [89] tackled source-free domain adaptation challenges in visual emotion recognition with a fuzzy-aware loss, effectively mitigating noisy pseudo-label issues in the target domain. Collectively, these advances reinforce the necessity of robust, noise-tolerant adaptation methods in diverse affective computing scenarios.
We employ a feature-based domain adaptation approach, focusing directly on visual modalities to preserve the integrity of emotional data, thereby avoiding the information loss associated with methods like those using image captions for domain adaptation [36]. This approach is exemplified by the FaceNet architecture, which learns mappings from facial images into representational spaces based on face-similarity measures [65]. However, these FER models often falter when applied to non-facial images, necessitating adaptations for broader IER applications. Our system adapts pre-trained FER models to effectively recognize emotions in generic non-facial images using this feature-based domain adaptation strategy. Employing discrepancy loss to align the feature distributions of FER and IER models without altering the network architecture ensures optimal transfer and adaptation of learned features across domains [63]. This method addresses the challenge of varied emotional cues in different image contexts and is capable of handling diverse datasets. It not only enhances the accuracy and applicability of our system but also interprets a wide range of visual features, significantly improving the robustness of emotion recognition.
@@ 2.3 Interpretability and Explainability
Interpretability and Explainability are increasingly vital in the field of AI and machine learning, particularly within the domain of emotion recognition [37, 35]. These concepts are essential for building trust in AI systems, as they allow users to understand and validate the decision-making processes of the models [7]. Interpretable emotion analysis has become a significant trend, with research expanding across both visual [72, 57] and other modalities [42, 33], underscoring the need for systems that users can understand and interact with confidently.
@@
With the increasing complexity of emotion recognition systems, the need for clarity in how decisions are made is crucial. Recent efforts by Kang et al. [24] have introduced a privacy-preserving adversarial domain adaptation approach, which not only protects sensitive information but also clarifies feature group processing, linking domain adaptation with clear, actionable insights. Asokan et al. [4] utilize concept activation vectors to illuminate the contributions of multimodal emotion cues, like facial expressions and vocal tones, to decision-making processes within models. Furthermore, Malik et al. [51] have enhanced the transparency of facial emotion recognition systems by localizing and accentuating critical facial regions, significantly boosting user trust and understanding of automated affective assessments.
Our research builds on these foundational works by proposing a system that not only explains emotion recognition outcomes but also provides insights into the training dynamics of network weights. This is crucial for the accurate interpretation of subtle emotional cues, which is essential in effective human-computer interaction. By enhancing transparency and interpretability, we aim to create more reliable AI tools for emotion analysis [34]. Our findings show that integrating real-time data interpretation and dynamic adjustment of network weights not only meets but exceeds theoretical expectations and outperforms traditional models in initial tests. These results underline the practical benefits of our approach in effectively addressing gaps in existing models.
@@ 3.3 Image Emotion Recognition
Preprocessing: Prior to training, significant preprocessing was applied to both datasets to ensure consistency and reliability of the input data. Initially, all images were resized to a standard dimension of 224x224 pixels and normalized to have zero mean and unit variance to mitigate any model biases due to variations in image scale and illumination. To address discrepancies in labeling conventions across datasets, a mapping scheme was devised that aligns the emotion labels from different sources into a unified set of categories. Furthermore, given the inherent class imbalance in the datasets, a combination of oversampling the minority classes and undersampling the majority classes was employed to ensure an equitable representation of all classes in the training process. These preprocessing steps are crucial as they significantly influence the model’s ability to generalize across diverse inputs and directly impact the performance outcomes discussed in Section 4.4.
Domain Adaptation Strategy: We employ a feature-based domain adaptation strategy using the same network architecture for both FER and IER tasks. This approach utilizes discrepancy loss to effectively align the target domain’s distribution with that of the source. The discrepancy loss, which is minimized after each training epoch, helps the adapted IER model accurately learn the IER domain’s distribution, despite the differing marginal distributions of data points between the FER and IER domains. This alignment is critical for the model’s ability to perform consistently across both domains. The overall loss function is detailed in Eq. 1.
@@
Theoretical Underpinnings of Discrepancy Loss: Discrepancy loss plays a critical role in our domain adaptation framework by quantitatively assessing and minimizing the disparity between feature distributions of the source (FER) and target (IER) domains. It measures the divergence between the probability distributions of features extracted from the FER and IER domains, specifically at the last hidden layer of the network. As incorporated in the overall loss function (Eq 1), $\mathcal{L}_{\text{discrepancy}}$ is instrumental in aligning these domains. The alignment facilitated by discrepancy loss is not merely a statistical alignment of distributions but a strategic enforcement of feature space homogenization. This process reduces the model’s sensitivity to the domain-specific nuances of emotional expression, which are prevalent when transitioning from FER to IER. By minimizing the discrepancy loss, we effectively guide the model to find a common feature space, where both domains’ statistical properties are not just aligned but are functionally indistinguishable in terms of their contribution to the model’s output. This ensures that the learned representations are robust, maintaining high accuracy and consistency across varied emotional datasets and contexts. Moreover, the use of discrepancy loss aids in overcoming overfitting to the domain-specific traits of the training data by promoting domain-invariant feature learning. This is critical for applications in real-world scenarios where the diversity of emotional expressions and the contexts in which they are presented can vary significantly from the conditions seen during training. By ensuring that our model learns to generalize across these domains effectively, we enhance its usability and reliability in practical deployments where domain variability is the norm.
Impact of Discrepancy Loss Function: The role of the discrepancy loss in our domain adaptation strategy is multifaceted. It is instrumental in training the model to reduce the divergence in output probabilities between the FER and IER domains for similar emotional inputs, facilitating a smoother transition from FER to IER tasks. Additionally, it allows the model to generalize more effectively across varied image contexts where emotional expressions differ significantly from the training data. This enhanced adaptability is crucial for deploying the model in real-world applications where it must perform reliably across different scenarios and datasets.
@@ 3.4.2 Layer-wise Interpretability
3.4.3 Comparison of Interpretability Approaches
The DnCShap technique compares with other interpretability techniques such as SHAP [49], LIME [62], and Grad-CAM [66] as follows.
Computational Efficiency: DnCShap enhances computational efficiency by incorporating a Divide & Conquer strategy to compute Shapley values, reducing the time complexity from exponential, as is typical with traditional SHAP, to linear. Our method computes approximate Shapley values in linear time, compared to the quadratic or higher time complexity of SHAP and LIME.
Interpretability Depth: Unlike Grad-CAM, which provides a coarse localization map indicative of relevant regions in the image, DnCShap offers detailed attributions at a pixel level, enabling finer interpretations of the contributions of specific image features to the emotion recognition task. This depth is particularly crucial in understanding subtle emotional cues in facial expressions.
Applicability to FER: The ability of DnCShap to provide detailed and computationally efficient interpretations makes it particularly suited for real-time FER systems, where understanding the decision-making process of the model is as important as its accuracy.
Table 1: Comparison of Interpretability Approaches.
Method Computational Complexity Detail of Interpretation Applicability to FER
SHAP High (Exponential) High (Pixel-level) Suitable
LIME Variable (High) Medium (Segment-level) Suitable
Grad-CAM Low (Constant) Low (Region-level) Limited
DnCShap Low (Linear) High (Pixel-level) Highly Suitable
The observations, presented in Table 1, demonstrate that DnCShap not only outperforms existing methods in terms of computational efficiency but also provides deeper insights, making it an ideal choice for advanced FER applications where both speed and interpretability are crucial.
Computational Efficiency and Scalability: DnCShap enhances computational efficiency by incorporating a Divide & Conquer strategy, which significantly reduces the computational overhead associated with calculating Shapley values. Typically, computing exact Shapley values in traditional SHAP has exponential time complexity. In contrast, our method manages to approximate these values in linear time, which is particularly advantageous when scaling to more complex models or larger datasets. To demonstrate the scalability and runtime efficiency, we conducted benchmark tests across datasets of varying sizes and complexity, which show that DnCShap maintains efficient performance without compromising interpretability, even as dataset size increases. These tests help substantiate DnCShap’s practical utility in real-world applications where both computational resources and time are limiting factors.
@@ 4.1 Datasets
The datasets utilized in this study are characterized by their large scale and diversity, encompassing a broad range of emotions and scenarios. These qualities make them exceptionally well-suited for the demands of both facial and image emotion recognition.
FER Datasets: The FER datasets used in this study are widely recognized in the research community for their robust and varied annotations of facial expressions, which are critical for the development and evaluation of FER systems.
  • FER13 [9]: Includes 35,887 emotion-labeled images, utilized extensively for its diversity in facial expressions and occlusions, facilitating robust model training.
  • FERG [3]: Comprises 55,767 annotated images known for their animated representations, which help in understanding the subtleties of facial expressions in animated contexts.
  • JAFFE [23]: Contains 213 images, often used for its high-quality depictions of basic emotional expressions by Japanese female models.
  • CK+ [48]: Features 927 images with labeled emotions, widely used for its sequential imaging that captures the onset and apex of facial expressions.
IER Datasets: These datasets are crucial for advancing research in image emotion recognition by providing a diverse range of contexts and settings.
  • IAPSa dataset [52]: Comprising 395 images, this dataset is pivotal for its role in psychophysiological studies and is frequently used to validate emotion recognition models due to its reliable emotional annotations.
  • ArtPhoto dataset [50]: Contains 806 art photos labeled with emotions by the artists, offering a unique dataset that bridges art and emotion, providing insights into the aesthetic elements that evoke emotional responses.
  • Flicker & Instagram (FI) dataset [82]: With 23,308 images, it represents the largest dataset used in our study, showcasing a wide variety of real-world scenarios that are instrumental in developing robust IER systems.
  • EMOTIC dataset [28]: Features a comprehensive collection of 23,571 images in diverse contexts, labeled with both continuous dimensions and discrete categories, making it invaluable for understanding emotional expressions in naturalistic settings.
4.2 Emotion Class Maping and Hyperparameter Setting
4.2.1 Emotion Class Mapping
In this work, we consolidate all emotion annotations to four discrete classes, i.e., happy, sad, hate, and anger. We chose these four because they are the most frequently used categories in state-of-the-art Image Emotion Recognition (IER) studies involving the same benchmark datasets: IAPSa, ArtPhoto, Flicker&Instagram (FI), and EMOTIC. As discussed in Section 4.4.2, many prior methods on these datasets also focus on these four dominant categories for consistency and reliable sample sizes.
Rationale for Merging: Although some datasets originally contain up to seven labeled emotions (e.g., amusement, contentment, disgust, fear, surprise, etc.), not all of these are well-represented or consistently annotated across the four datasets. For instance, IAPSa and ArtPhoto include additional labels like amusement or contentment, but each has relatively few samples. Likewise, disgust is labeled in some, but not all. To avoid excessive class imbalance and enhance cross-dataset alignment, we adopt the following merges aligned with Plutchik’s emotion wheel [59]: contentment and amusement are merged to happy whereas disgust is merged to hate. In this manner, each of the four datasets has sufficient examples in the four selected emotion categories, facilitating robust training and evaluation.
Impact on Model Performance: By consolidating underrepresented labels into these four commonly used categories, we mitigate the risk of overfitting to sparsely populated classes and reduce confusion between closely related affective states (e.g., disgust vs. hate). Empirically, we observe more stable training curves and higher overall accuracy compared to attempts at using six or seven fine-grained categories. In particular, merging disgust into hate and grouping contentment/amusement with happy provides clearer class boundaries, resulting in better cross-dataset generalization. While we focus on these four discrete classes for the present work, the underlying domain adaptation framework could be extended to additional discrete labels if future datasets offer consistent and balanced annotations.
4.2.2 Training Strategy and Hyperparameter Setting